Please help..
solve:
1)4x + 3y =7
2) 16x + 12y = 28

Thank you so much....Rebecca

Found 3 solutions by JBarnum, nyc_function, jim_thompson5910:
Answer by JBarnum(2146)

when u multiply the top equation by 4 then subtract the eqations u get 0=-21
so x=y
so if u look at the first equation theres only 1 possible number that x and y can both be and same for 2nd equation
4+3=7
16+12=28
so
x=1
y=1

Answer by nyc_function(2741)

HINT: Use the substitution method for systems of linear equations in two variables. Can you do it now?

Answer by jim_thompson5910(35256)

Hint: If you multiply both sides of the first equation by -4, you get -16x-12y=-28. Now add this to the second equation and you'll find that the 'y' terms cancel out. This will leave you with an equation in which you can solve for 'x'.